Mary Arnold – Faith and the Everyday Practice of Science

Mary, this is a very good start on an essay concerning Grinnell’s ideas about complementarity of science and religion. However, I don’t think you have really “nailed” a description of the differences between the two fields, at least according to common thinking (what Grinnell is arguing against). Your first paragraph points to the distinction between pragmatic concerns and theoretical concerns. But that is not so clear-cut (see comment 1 below). Your second paragraph says the fundamental conflict between science and religion is in how they draw conclusions. There certainly is a difference there, but is that the most central difference? It might be good to start a revised version of this by addressing another question: What are the ultimate goals of science and religion, and how are they different (and/or similar)?

It is also important to keep a bit of critical distance and ask the question whether Grinnell’s suggested reconciliation of religion and science is really possible.

Specific Comments

1. It sounds as though you are equating science with the pragmatic and religious with the theoretical here. But I would say that science and mathematics certainly have a theoretical side as well as a pragmatic side. So the difference between science and religion is not only the difference between pragmatic and theoretical concerns.

2. Everyday Practice of Science is a rather unclassifiable book, so it could be called “novel,” meaning new or unusual. I’m guessing that is what you were thinking here. But when I read this the first time, it seemed that you were calling the book a work of fiction. It would be better to rewrite this to eliminate the possible confusion.

3. I think it would be better to say here that courts held that the introduction of creationist science curricula was a breach of the separation of church and state. This is a somewhat controversial legal issue. Your flat statement makes it sound too cut and dried. Also, it’s interesting to consider the way creationism is often presented as a “scientific” theory and an alternative to Darwinian evolution. Why do you suppose the supporters of creationism try to do that?

4. This is still a part of an expression of Grinnell’s opinion. It would be good to make that clearer because there are certainly lots of scientists who would disagree violently with this statement.

5. I am not sure what you meant by this. Do you mean science cannot explain the human feelings of a connection to the divine??

6. This idea that embracing complementarity could lead to mutual respect and acceptance between religious and scientific viewpoints is a very good one. You could say more about that! But on the other hand, is Grinnell’s proposal realistic? as a way to reconcile science and religion?

Amanda Cabrera – One Natural World, Two Visions

I got your email message accompanying the paper, and I agree that this is still somewhat rough and unfinished. I hope you will have the time to think over the following comments and resubmit a revised version of the paper. It is also important to keep a bit of critical distance and ask the question whether Grinnell’s suggested reconciliation of religion and science is really possible.

Specific Comments

1. Saying “integrated” here sounds a little strange. Many scientists might say that when religious concerns impact how science is taught and practiced, then it is more an intrusion than an integration. (The primary meaning of integration is inclusion or mixing in a more neutral way.) Your second sentence is making a very different point (following Grinnell). So this probably deserves to be expanded and explained a bit more – you need to give the reader a clue that you will be dealing with two very different things.

2. The people who have been agitating for “creationist science” in schools are primarily fundamentalist Protestants. In recent times at least, the Catholic Church has been less prone to that sort of literal interpretation of Scripture and much more open to modern geological theories of the age of the Earth and biological theories of evolution. So I think this sentence is somewhat inaccurate. (By the way, I’m thinking of the Church as the whole community of Catholic believers and Catholic institutions – colleges and universities included – not just the Church hierarchy.)

3. Grinnell’s idea of “intelligible design” is also a very good point to spend more time on. The idea here is the feeling (which comes down to a form of faith, for Grinnell at least) that there are underlying patterns in Nature that humans are able to understand.

4. I don’t really see the connection between what you are saying about Biblical literalism and Grinnell’s complementarity here. This is something you probably want to rethink.

5. Is it faith that is expressed through the reasoning of where to go on the bicycle? Isn’t this sort of reasoning more closely related to ethics or morality? Aren’t those different from faith? (That is, religion also includes ethical and moral statements of what we should do.)

6. Could you say more about why you think Grinnell’s dabbling in philosophy is dangerous here? Are you saying he is not qualified to offer those opinions because he is “only” a scientist?

Stephanie Craig – The Battle Between Effect and Passion

I like your first draft paper very much – it has a number of good points and it strikes a good balance between your personal experience and the ideas expressed by Grinnell in his book. There are a few places where you had typos or minor slips. Also there were a few examples of what I would call overly colloquial language – “zapping the life out of” (page 1) in particular.

Specific Comments

1. I gather that a lot of students feel the way you describe here (that is, that having to explain a solution takes the fun of out finding the solution). Could you say more about why you have that feeling? I don’t really understand that myself, since I always think that the best way to make sure you really understand something is to try to explain it to someone else. Any gaps in one’s thinking become painfully clear then, either because of the uncomprehending stares that you get from one’s audience, or from the feeling of “hand-waving” that comes from trying to dance around a point that is not entirely nailed down.

2. I like this example very much – I think it is exactly right!

3. You are also correct that statistics is often used in the way you describe – as a way to summarize a large collection of information and make it easier to deal with. You seem to be saying that this is like the process of writing a short and to the point research article without any of the distracting extra information that might have been developed along the way to the final results. Am I catching your point here, or did you mean something different? Also, aren’t there times when you might discover new things in the process (patterns that were not apparent in the whole data set at the start). That is, couldn’t working with the statistics be part of the discovery process, in addition to or rather than the credibility process?

Josh Hauser – Practice of Science and Mathematics

Your first paragraph was a little rough (see the suggested corrections marked on the paper). But after you got started, your writing improved quite a bit. One suggestion – “kids” sounds too informal for this kind of writing. What you really mean is “students” so say it that way.

I agree with a lot of what you are saying here, and you present your points pretty well.

Specific Comments

1. This is a very important point. This kind of training in logical thinking and mental discipline is also one of the traditional reasons for including mathematics in school curricula. This goes all the way back to the ancient Greeks.

2. I’m not sure it’s really accurate to say that there is “much room for extra information” in research articles. After all, textbooks are usually hundreds of pages long, but articles are only rarely longer than 20 or 30 pages. There is a difference too in the level of knowledge assumed of the readers. In research articles, the audience is the community of experts in a field, so not all of the background needs to be filled in.

3. I’m not sure I follow what you are saying here. It is not that common, I think, for later articles to re-work the data from earlier ones. So I don’t think “taking the results from a previous experiment in order to find different patterns and characteristics” happens that often (at least not in the parts of science I know the most about). Instead, people might try to reproduce other scientists’ results by redoing the experiments in a different way and generating new data of their own.

Paige Carr – Faith in Religion and Science of the Natural World

Your essay is well-structured and presents a good synthesis of Grinnell’s ideas about the relation between faith, science and religion. One suggestion: It is also important to keep a bit of critical distance and ask the question whether Grinnell’s suggested reconciliation of religion and science is really possible. Other than that, I only have a few small corrections and comments.

Specific Comments

1. You say that “without religion, there would be no point in discovering the reason why things in life work.” Does Grinnell say anything like that? Or is this a personal opinion of your own? It would be good to say more about this and try to justify it, because I think this is a statement that many people might disagree with. Isn’t it possible for an atheist to be a scientist and want to understand how nature works? Can’t an atheist be a good, moral person and be concerned with the welfare of others? Is religion the only source of moral and ethical values that provide the basis for a good life?

2. Toward the end of the essay when you are discussing Grinnell’s idea of complementarity, there are a few instances where you are misusing the word. This is a noun – the property of being complementary. The adjective form is complementary.

Nick Foster – (no title)

Your view of three stages in the process of the development of scientific truths – the notebook stage, the article stage, and the textbook stage – has a lot of truth to it. Still, I think you could say more about why the later stages in the process look so different from the notebook stage. See the comments below for some ideas. Your writing is generally very good, although there are a few cases where I think your spell-checker and/or your proof-reading may have let you down(!) Also, notebook and textbook are single words.

Specific Comments

1. I think it might be a slight overstatement to say that “usually” the final result of a scientific investigation “is in no way related to what [the scientist] originally intended to discover.” Grinnell was saying that this happens more often than one might think if one believes the textbook picture of the “scientific method.” But there are times too when scientists do succeed in finding what they are looking for!

2. I think it’s important to realize that bias can also come into the choice of questions that the scientist thinks are important or that he or she decides to investigate. For example, the meeting I was at two weeks ago was the national convention of a group called SACNAS (Society for the Advancement of Chicanos and Native Americans in Science). One of the plenary sessions was a pretty interesting discussion about the ways modern “evidence-based” medicine can learn things about treating Native American patients from traditional tribal healers. Those questions probably would not even occur to doctors who had no experience with Native American cultures.

3. Careful – a scientist’s notebooks are almost never published. They are retained to provide back-up if questions arise concerning the results the scientists claims to have found. But usually they are not shared directly with other scientists.

4. Here’s a question to think about. Are all the misadventures and false starts that a scientist may have had on the way to discovery even relevant to the credibility process? Might they even be a distraction from the real issue of determining whether the results are credible? Could that explain the difference in the ways the different stages are presented?

5. This sentence does not read correctly. I think you might have left out a part of what you meant to say, although what you intended is pretty clear from the context. What does this say about the textbook format as a learning tool, though? If better-quality learning happens through discovery, shouldn’t teachers try to include more of that kind of experience in students’ education?

Brandon Nunn – Frederick Grinnell’s Everyday Practice of Science

This is, for the most part, a very good essay on Grinnell’s ideas about how scientific practice differs from the textbook picture of the “scientific method.” The one criticism I have is that the last page or so seems to just rehash points you have made previously. You don’t bring things to a strong conclusion. There are some things you could add or expand earlier in the paper, though, so the length should not be an issue if you want to do a revision. See the specific comments below for some ideas on this.

Specific Comments

1. Actually, in mathematics, the proof of a statement is supposed to be a way to convince others that the statement is true. If you saw the proofs (and not just worked out examples, procedures to follow, and so forth) then your teachers probably thought that they were showing you why those facts were true(!)

2. I think there are really two points here. One is that most experiments just don’t work at all. The lab animals die, or the meter doesn’t work, or the computer was misprogrammed, etc. Then sometimes, there are unexpected results (the “serendipity” idea).

3. This high school chemistry course sounds like it was a great experience. You must have had a very dedicated, knowledgable, and confident teacher!

4. I think Grinnell would say that this effect is even more pronounced on the discovery side. Scientists don’t want to make extravagant claims that are later proved to be false because that would hurt their reputations even more.

5. You introduce the idea of intersubjectivity here without really explaining what Grinnell means by this term. This could be expanded quite a bit. Don’t assume the reader knows what Grinnell has said about this.

6. I agree that statistics could be used to discover patterns. Couldn’t those techniques also be used to demonstrate to others that a pattern observed more intuitively are “really there?”

Patrick Jarvis – Faith vs. Science

To me, your paper sounds almost angry in places and from your writing it seems that this topic has a large component of emotional resonance for you. It is sometimes difficult to write about topics like that because the temptation is to get carried away and overstate your ideas. However, I think the picture is really quite a bit less clear-cut than you seem to think. There are lots of “gray areas” and distinctions that need to be taken into account to get a balanced view of the issues that you are raising. Although you mention several points that Grinell makes about the relation of faith and science, you should see that he comes to a quite different final conclusion in his final chapter. I hope you will consider revising this paper. If you do, it might be good to refocus your topic and address Grinnell’s ideas in more detail. It is also important to keep a bit of critical distance and ask the question whether Grinnell’s suggested reconciliation of religion and science is really possible.

Specific Comments

1. In recent times (say the last 100 years), the situation is not at all as clearcut as you are indicating. Because the Catholic Church is such a large and diverse institution, it is really hard to pin down a single “Catholic” point of view on topics like this. (By the way, I’m thinking of the Church as the whole community of Catholic believers and Catholic institutions – colleges and universities included – not just the Church hierarchy.) Within that group, I think it’s accurate to say that literal interpretation of Biblical scriptures about the age of the Earth, and rejection of the idea of evolution based on those interpretations, is far less common than in some more fundamentalist Protestant groups. The Church positions against abortion, embryonic stem cell research, and so forth are based not only on passages from the Bible, but on a whole moral tradition of respect for human life in every form. Biblical literalism has not really been a part of the wider Catholic tradition for some time.

2. I don’t think it is accurate to implicate the Catholic Church in the Scopes trial. Don’t forget that this trial dealt with teaching of evolution in a public school in the state of Tennessee. The issue was whether a state law could forbid teaching that topic in biology classes (based on a literalist, Protestant interpretation of the Bible). While some Catholics certainly held opinions one way or the other, this trial did not involve the Church directly.

3. The charge that humans are trying to “play God” by engaging in research about stem cells could also be a scare tactic to stir up emotion against a legitimate scientific object of study.

4. I really doubt that bringing Dan Brown’s novels into this discussion adds anything to your arguments. As you say, the plots are farfetched and sensationalized. While they are admittedly fun to read, the picture they paint of the Catholic Church is hardly a realistic or realiable one.

5. If you choose to revise the paper (and I hope you will), it would be really good to “dig into” Grinnell’s idea of complementarity of science and religion in more detail. Personally I think that his ideas are one way to reconcile the conflicts between science and religion that you have highlighted, at least for an individual scientist.

Carl Losito – Faith in Science and Religion

Your essay makes a number of very good points, and you have raised an important possible objection to Grinnell’s vision of “peaceful coexistence” of science and religion in the notion of complementarity. So, the main comments I have are not so much about what you say, but rather about how you say it. I think there are a number of places that could stand some rewriting to simplify and clarify what you wrote the first time. I have tried to point out the simpler ones where just changing a word or two would make a sentence read more clearly directly in the paper. The more extensive comments are below:

1. I didn’t quite follow this transition between the idea of Nature changing over time and the idea of researchers studying the changes to see if new antibiotics are possible. The transition is too abrupt for a reader to follow your thoughts. I think you need a sentence or two more here to say first that science looks for patterns in that changing landscape. Then, add that some of the patterns they find can have practical applications such as chemicals with antibiotic properties produced by natural organisms.

2. It’s true that nature can change over time. An important part of what Grinnell is saying, though, is that however the natural world develops, scientists still must have a sort of faith that the patterns that are present are things that they (the scientists) are capable of finding and understanding – that there is an “intelligible design” (Grinnell’s term) underlying the natural world.

3. I’m not sure what this phrase “more difficult to welcome” means. Could you explain this more fully?

4. Instead of “does not imagine,” wouldn’t it be more accurate to say something like “does not believe that a person should have that sort of power over another person.”

5. This sentence was another “tough nut” for me. I don’t understand what you are trying to say.

6. This point could be expanded and explained more. Science and technology provide the tools to do things; religion furnishes (among other things) the moral compass that leads to good choices about how the technology will be used.

7. This is a really important point. You are asking, in effect, how open fanatically religious people would be to the complementarity and dialogue that Grinnell sees as the ideal relation between science and religion. Is Grinnell being too optimistic about the chances for these ways of looking at the world to coexist?

Shelby Proft – The Practice of Science

You have some good things to say here, especially about the different ways statistics might be used at all stages of the process of discovering and establishing the credibility of a scientic result. However, the way you present your ideas sometimes makes it very difficult for a reader to decipher what it is you are trying to say. Your writing is sometimes unclear and overly complicated. Good writing does not have to make use of long, convoluted sentences and “vocabulary list” words. My advice to any “apprentice writer” (a class that includes almost all college first-year students!) is that simpler, more direct ways of saying things can be even more effective. Striving for simplicity can also help you refine your ideas and come to a clearer understanding of just what it is that you are trying to say(!)

I “marked up” your essay especially heavily with suggested rewordings and reorganizations of some sentences. I also indicated a number of possible deletions of phrases or sentences that are unnecessary or repetitive. There are also some more extensive specific comments below. Please take a close look at all of these and come talk to me in office hours if they are not clear or if you want to discuss what I am suggesting.

Specific Comments

1. This first sentence is not even especially complicated, but I do not know what it means. What does “measured down to” mean here? The first sentence of an essay has to be especially clear to give the reader the introduction to what you are going to say.

2. In this sentence, I don’t know what you mean by “projects.” Are you saying that the typical science fair project that a student does gives a false impression of what science is actually about? Also note the possible confusion for a reader – when I see “science fair projects” I almost automatically think of “projects” as the noun referring to the activities the students carry out and present at the fair. Be alert to what the reader will be getting out of what you write.

3. “The lack of clarity ... is clear.” Did you mean to say this? There are places for jokes or puns in formal writing, but I don’t think this is one of them.

4. Your last paragraph is not a very strong conclusion to your paper. The final sentence just meanders without getting anywhere. The conclusion of an essay also requires special attention. The idea is to tie things up for the reader and reiterate the most important point you want her or him to take away from what you have said.

Andrew Coury – Identifying the Ideal Way of Practicing Science

I do not think your title really describes what you end up saying in the paper. But the essay itself is generally very good. Your writing is, for the most part, clear and direct. The main comment I have (see below) is that you have essentially made two different claims about where statistics “fits into the picture.” While they are not entirely contradictory, they do conflict to a certain extent. You should think this through again and decide which of the two positions you want to assert. Of course it may be the case, too, that you ultimately say you think statistics fits both in the “discovery side” and the “credibility side” of the picture(!)

1. Your first sentence involves a sort of pun because the word “practice” has two different meanings. Grinnell is using the word in the same sense it has in the phrase “common practice” – that is, practice as “the way things are done.” The saying, of course, involves “practice” in a different sense – the idea of preparing for doing something by repeating exercises designed to develop the skills that will be needed (think band practice, or football practice). I’m not sure that “practice makes perfect” really has that much to do with what Grinnell is saying, although you could argue that it does, I suppose. If you want to make a connection, you would have to say more.

2. It would be a good idea to put “Professor Anybody” and “Professor Particular” in quotes and identify them as coming from Grinnell’s discussion of how science is done.

3. A main reason that traditional geometry courses have been included in school curriculums has always been that the process of developing and writing proofs is thought to be good training in logical thinking and also good training in the careful presentation of results. Those skills are intended to be transferable to other areas besides geometry. “Practice makes perfect” in action!

4. You say “research articles should provide more detail ... ” Why? Is that detail needed for their intended readers? Is that added detail about how the results were discovered even relevant to the “credibility” process? (Also, please use the format given in the assignment for identifying quotations from Grinnell.)

5. What does “adversity in validating thoughts” mean? Are you trying to say that research papers have to undergo a more stringent process of examination before they achieve a degree of credibility?

6. (and 7.) This is what I meant above when I said you are arguing both sides of the question about where statistics fits into the picture. Your statements are not exactly contradictory but they are close to contradictory. Think this through again and decide which it is (or whether, in fact, it is both).

Kyle Walraven – Faith in Religion and Science

This is an excellent, well-argued, and rather persuasive essay. You took a somewhat unusual approach by criticizing Grinnell’s point of view, but I think your reservations about Grinnell’s ideas are generally well-founded.

If I were trying to defend Grinnell here, I would say that his picture of complementarity might work best as a way for an individual scientist (like Grinnell himself) to reconcile his or her profession as a scientist with the religious impulses that he or she might also have. (Think of the “bicycle and rider” metaphor – science gives the tools for doing things, religion gives the guidance for what things should be done.) Grinnell also seems to be thinking on a very abstract level – he seems to be thinking about the idea of religion, for instance, rather than the specifics of any particular religion. But if you think on a wider, or a more realistic, level (whole groups of people or institutions, actual particular religions), the conflicts you point out definitely tend to work against the coexistence that Grinnell is hoping for.

Your writing is generally strong and clear. I have marked a few minor suggestions on the paper.

Specific Comments

1. You might want to put in a separate sentence here to give an example to replace the phrase “such as the sun” that I deleted to make this sentence flow better. Also, this question of why religions of the past have died out is an interesting one. Why do you think that has happened? What replaced the Greek (and ancient Roman) religions?

2. Even the first statement here is questionable. Is fundamentalist Islam (or fundamentalist Christianity, for that matter) willing to accept all of science? parts of science?

John Vatalaro – (no title)

Your essay is a good summary of Grinnell’s ideas about faith and science. But I think it is also important to keep a bit of critical distance and ask the question whether Grinnell’s suggested reconciliation of religion and science is really possible. Your writing is also generally good (direct and clear). There are a few somewhat minor corrections that I marked on the paper itself. You might also want to think about the following comments.

Specific Comments

1. Isn’t this difference between religion and science a really big obstacle to the kind of coexistence that Grinnell is proposing? Is Grinnell underestimating the difficulty of this coexistence by saying “religions change too?”

2. While some traditions of interpretation of Jewish scripture have some of the scientific aspects that you are describing, I’m not sure that this is a completely representative example of the way religions operate. Think of fundamentalist Protestant groups where the Bible is taken completely literally. (That is, for instance, the universe was created in 6 literal days, about 4000 years ago.)

3. Why do religion and science “have to be synchronized?” There are some very outspoken atheist scientists who argue that there is no place at all for religion in a rational view of the world. (Look up the biologist Richard Dawkins for a really extreme example.)

4. Is the purpose of religion really to “solve problems?” Or is that what you are saying here?

Dan Schmader – Faith in Science and Religion

Your essay is a pretty good summary of some of Grinnell’s ideas about faith and science. You seem to be focusing on the level of individual experience, and I guess you could say Grinnell does the same thing. But I think it is also important to keep a bit of critical distance and ask the question whether Grinnell’s suggested reconciliation of religion and science is really even possible on the level of groups of individuals or religious and scientific institutions. Can the differences really be reconciled? Your writing is mostly good, although there are a few unclear places that I tried to mark on the paper or point out in the following comments.

Specific Comments

1. Check the spelling of Grinnell (!)

2. That is perhaps a part of what Grinnell is saying, but there are other aspects of this as well. It would be good to look over his comments on the need for scientists to have a kind of faith that Nature has patterns that human beings can understand – the idea of “intelligible design.” That is another important aspect of what he is saying in this connection. What you say later on page 2 is also getting at this idea but you never really state it explicitly.

3. Although this is somewhat unclear from the historical record, most people think Galileo was mistreated to a much greater degree than just being denied credit for what he had discovered. He was apparently forced to renounce his ideas about the Earth moving around the Sun and forbidden to discuss them in public by the Church.

4. This is the place I indicated above in comment 2. There is more to what Grinnell is saying about the role of a sort of faith for scientists than you are getting at.

5. I am not sure what you mean by this. Does the “it” refer to the self? I think a lot of people would say religion is a lot more than just a way to self-understanding.

6. This is another place where what you are saying is not so clear. Do you mean that the student was asking Grinnell to justify why he existed? If so, I think there is an even more basic point that comes up in a lot of first-year philosophy classes. Namely, can any of us really say that we know we exist in a real world outside of our own minds? How do we know what we think we know about that world? In other words, this is a question of epistemology – the branch of philosophy that deals with the foundations of knowledge and addresses questions such as “how do we know what we know?”

Greg Monfette – Faith in Science

Your writing here is generally good. However, I think there are some aspects of what Grinnell is saying about faith in science, and the relation of science and religion, that you have not completely understood. It would be good to go back over what he is saying and try to address some of the points in the specific comments below.

Specific Comments

1. “Adhesion” is the wrong word here, I think. You want “adherence to” or maybe even better “agreement with.”

2. I think this example is one where it is almost too easy to reconcile relgious ideas with scientific theories, because they are already very close. The scientific Big Bang theory omits any speculation about what caused the Big Bang to occur or why it happened. Introducing a Creator is one possible explanation for what might have caused it to occur or why it happened. Since science doesn’t try to answer that question, there is no real conflict. In other words, there is not that much of a disagreement. It is different, I think, for other areas where modern scientific theories have a greater degree of conflict with traditional religions. (For instance, how could you reconcile a literal interpretation of the Bible, such as is common in some fundamentalist Protestant sects, with scientific estimates that the Earth is between 4 and 5 billion years old?)

3. This seems too general to me. There is a lot of variation within religions and between religions. For example, the people who have been agitating for teaching “creationist science” in schools are primarily fundamentalist Protestants. In recent times at least, the Catholic Church has been less prone to that sort of literal interpretation of Scripture and much more open to modern geological theories of the age of the Earth and biological theories of evolution.

4. Grinnell’s complementarity is much more subtle than you are saying it is here. It is not only the idea that people’s views or the teachings of a religion can change. It has more to do with the relation between the scientific and religious viewpoints. This is one aspect of the paper that should be rethought and reworked.

5. I think it is also important keep a bit of critical distance and ask the question whether Grinnell’s suggested reconciliation of religion and science is really even possible on the level of groups of individuals or religious and scientific institutions. Can the differences really be reconciled?

Jack Karlovec – Practice of Science

This a start but it is not a good effort. I hope you will take the following comments seriously and rethink and rework this essay.

First, there are a number of very basic spelling and grammatical errors here that must be corrected. The most glaring ones: please make sure you understand the difference between “then” (sequence in time) and “than” (comparison). Also make sure you understand the difference between “its” (possessive) and “it’s” (contraction for “it is.”) In formal writing, using the contractions is generally frowned on (unless you are quoting direct speech). So if in doubt, spell it out. If you are not sure about any of this, ask me.

Even worse than the technical problems, the essay really does not say very much and it does not show a good understanding of what Grinnell was getting at. In addition, much of it repeats the same ideas over and over without any development. Did you try to outline what you wanted to say before you plunged in to write, or did you just grind it out? It reads as though you did the latter. I marked some corrections and suggestions about rewordings on the paper itself. Please think about the following comments before you start rewriting.

Specific Comments

1. I do not know what you mean by saying that scientists and mathematicians are “more important” on the discovery side. Everything that Grinnell is discussing deals with what scientists and mathematicians do.

2. This is overstated. I think it is better to say that Grinnell points out that failure is sometimes a part of scientific work. That failure might sometimes be worthwhile if it leads the scientist to a different and ultimately successful way of approaching the problem. But that does not always happen and the experience is not always worthwhile.

3. This is one of the places where you have really just repeated the same statements several times without developing them or drawing any conclusions.

4. You seem to be focusing on the fact that textbooks can seem tedious (maybe especially so for readers who are not especially interested in the subject). That is certainly true. But what Grinnell is saying is partly that he thinks that the situation might be different if textbooks showed more of what the scientists actually did to find the facts that those textbooks present. Did you think about that? Do you agree with it? Can you see ways that might make calculus books (to take just one example) more interesting? On the other hand, might there be good reasons why textbooks are written the way they are?

5. I really do not agree with this statement that anyone can write a textbook (if that is what you meant to say). Writing a good textbook is different from doing research. But it is also really hard and requires knowledge both of the subject and of effective ways to present it.

6. I think you could probably make just as good a case to argue that statistics are primarily important on the “credibility” side. There are times when scientists may have an intuition that a pattern is present without really being able to say why they think it is so. Doing the number-crunching to compute a correlation coefficient (for instance) can also be a way to convince other scientists that the pattern they think they see is really there.

Jonathan Ye – (no title)

This is an excellent and persuasive essay on some of the educational effects of the standard methods of presenting science and mathematics in textbooks and courses. I agree with most of what you are saying here, and I think that things have to change too. Your writing is generally very strong too. There are a few places where I think you chose the wrong word or where you might reword things for greater clarity. See the comments on the printout of the paper.

Specific Comments

1. I think what you are describing here is actually relatively rare. In mathematics (which is, of course, the literature I know best), there is a tradition of showing only the polished final result without any of the background or the false starts or setbacks that might have led the authors to the final conclusion. I think most scientific articles are pretty similar. This is Grinnell’s view too. The papers you read for your project sound very unusual.

2. It is possible that showing how something like the Pythagorean Theorem might be used in other contexts could make it seem more relevant and interesting to students learning geometry. But at the same time, doing that would require the students to learn some of the background needed to understand the applications. That would take time and also require some knowledge on the part of the teacher. Maybe you can see from these comments why this approach is not always followed(!)

3. This was one place where what you said was not entirely clear. Did you mean “definitions” or “proofs?” There are probably well over 200 different proofs of the Pythagorean Theorem that people have found and published over the years. Or did you mean that the concepts involved in the Pythagorean Theorem require lots of background (preliminary definitions) to understand?